# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 2.5. Pp-constructions
Definition 2.11 (fractional map).
Let $A$ and $B$ be sets. A fractional map from $B$ to $A$ is a probability distribution
$(A^{B},\mathcal{B}(A^{B}),\omega\colon\mathcal{B}(A^{B})\to[0,1]),$
that is, $\omega(A^{B})=1$ and $\omega$ is countably additive: if $S_{1},S_{2},\dots\in\mathcal{B}(A^{B})$ are disjoint, then
$\omega(\bigcup_{i\in{\mathbb{N}}}S_{i})=\sum_{i\in{\mathbb{N}}}\omega(S_{i}).$
@@
The set $[0,1]$ carries the topology inherited from the standard topology on ${\mathbb{R}}$. We also view ${}{\mathbb{R}}\cup\{\infty\}$ as a topological space with a basis of open sets given by all open intervals $(a,b)$ for $a,b\in{\mathbb{R}}$, $a<b$ and additionally all sets of the form $\{x\in{\mathbb{R}}\mid x>a\}\cup\{\infty\}$.
@@
Definition 2.12 (fractional homomorphism).
Let ${\mathfrak{A}}$ and ${\mathfrak{B}}$ be valued $\tau$-structures with domains $A$ and $B$, respectively. A fractional homomorphism from ${\mathfrak{B}}$ to ${\mathfrak{A}}$ is a  fractional map $\omega$ from $B$ to $A$ such that for every $R\in\tau$ of arity $k$ and every tuple $t\in B^{k}$ it holds for the random variable $X\colon A^{B}\rightarrow\mathbb{R}\cup\{\infty\}$ given by
$f\mapsto R^{{\mathfrak{A}}}(f(t))$
that $E_{\omega}[X]$ exists and that
$E_{\omega}[X]\leq R^{{\mathfrak{B}}}(t).$
We refer to [15] for a detailed introduction to fractional homomorphisms in full generality. If ${\mathfrak{A}}$ is a countable valued structure, we have the following handy expression for $E_{\omega}[X]$, which we sometimes use in the proofs:
$E_{\omega}[X]=\sum_{s\in A^{k}}R(t)\omega({\mathscr{S}}_{t,s}).$ (1)
All concrete fractional homomorphisms $\omega$ that appear in this paper are of a very special form, namely, there is a single $f\in A^{B}$ such that $\omega(\{f\})=1$. In this case, we also write $f$ instead of $\omega$. If $\omega$ is of this form, then for every $R\in\tau$ of arity $k$ and $t\in B^{k}$, the expected value in Definition 2.12 always exists and is equal to $R^{\mathfrak{A}}(f(t))$. If additionally ${\mathfrak{A}}$ and ${\mathfrak{B}}$ are crisp structures, then we call $f$homomorphism. It is easy to see that there are valued structures ${\mathfrak{A}}$ and ${\mathfrak{B}}$ with a fractional homomorphism from ${\mathfrak{B}}$ to ${\mathfrak{A}}$, but no fractional homomorphism from ${\mathfrak{B}}$ to ${\mathfrak{A}}$ of the form $f\in A^{B}$.
Lemma 2.13.
@@
Proof.
Suppose that there exists a fractional homomorphism $\omega$ from ${\mathfrak{B}}$ to ${\mathfrak{A}}$. Since $B$ is countable and $\operatorname{Aut}({\mathfrak{A}})\subseteq\operatorname{Aut}(\operatorname{Feas}({\mathfrak{A}}))$ is oligomorphic, it suffices to show that every finite substructure ${\mathfrak{F}}$ of $\operatorname{Feas}({\mathfrak{B}})$ has a homomorphism to $\operatorname{Feas}({\mathfrak{A}})$ (see, e.g., [3, Lemma 4.1.7]). Let $b_{1},\dots,b_{n}$ be the elements of ${\mathfrak{F}}$ and $b:=(b_{1},\dots,b_{n})$. By the (countable) additivity of probability distributions, there exists $a\in A^{n}$ such that $\omega(\mathscr{S}_{b,a})>0$. Let $f$ be the map that takes $b$ to $a$. Suppose that there exists $R\in\tau$ of arity $k$ and $t\in F^{k}$ such that $R^{{\mathfrak{A}}}(f(t))=\infty$. Since $\omega(\mathscr{S}_{b,a})>0$ and $\mathscr{S}_{b,a}\subseteq{\mathscr{S}}_{t,f(t)}$, we have $\omega({\mathscr{S}}_{t,f(t)})>0$, and thus $E_{\omega}[g\mapsto R^{{\mathfrak{A}}}(g(t))]=\infty$ by (1). Then $R^{{\mathfrak{B}}}(t)=\infty$, because $\omega$ is a fractional polymorphism. Hence, for every $R\in\tau$ of arity $k$ and $t\in F^{k}$ we have $R^{{\mathfrak{A}}}(f(t))<\infty$ whenever $R^{{\mathfrak{B}}}(t)<\infty$. Therefore, $f$ is a homomorphism from ${\mathfrak{F}}$ to $\operatorname{Feas}({\mathfrak{A}})$.
The final statement follows from the fact that every homomorphism is a fractional homomorphism and that ${\mathfrak{C}}=\operatorname{Feas}({\mathfrak{C}})$ for every crisp structure ${\mathfrak{C}}$. ∎
We say that two valued $\tau$-structures ${\mathfrak{A}}$ and ${\mathfrak{B}}$ are fractionally homomorphically equivalent if there exists a fractional homomorphisms from ${\mathfrak{A}}$ to ${\mathfrak{B}}$ and from ${\mathfrak{B}}$ to ${\mathfrak{A}}$. Clearly, fractional homomorphic equivalence is indeed an equivalence relation on valued structures of the same signature.
Definition 2.14 (pp-construction).
@@
Remark 2.15.
Let ${\mathfrak{A}}$ and ${\mathfrak{B}}$ be relational structures, and suppose that ${\mathfrak{A}}$ pp-constructs ${\mathfrak{B}}$, that is, there is a pp-power ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ which is fractionally homomorphically equivalent to ${\mathfrak{B}}$. It follows from Remark 2.6 that ${\mathfrak{A}}^{\prime}$ is essentially crisp and $\operatorname{Feas}({\mathfrak{A}}^{\prime})$ is a pp-power of ${\mathfrak{A}}$ with all relations primitively positively definable in ${\mathfrak{A}}$ when viewed over the domain $A$. By Lemma 2.13, $\operatorname{Feas}({\mathfrak{A}}^{\prime})$ is homomorphically equivalent to ${\mathfrak{B}}$. Hence, our definition of pp-constructability between two relational structures with oligomorphic automorphism groups coincides with the definition from [2].
The relation of pp-constructability is transitive: if ${\mathfrak{A}}$, ${\mathfrak{B}}$, and ${\mathfrak{C}}$ are valued structures such that ${\mathfrak{A}}$ pp-constructs ${\mathfrak{B}}$ and ${\mathfrak{B}}$ pp-constructs ${\mathfrak{C}}$, then ${\mathfrak{A}}$ pp-constructs ${\mathfrak{C}}$ [15, Lemma 5.14].
By $K_{3}$ we denote the complete graph on $3$ vertices. The following is a direct consequence of [15, Corollary 5.13] and [3, Corollary 6.7.13].
Lemma 2.16.
Let ${\mathfrak{A}}$ be a valued structure with an oligomorphic automorphism group. If ${\mathfrak{A}}$ pp-constructs $K_{3}$, then ${\mathfrak{A}}$ has a reduct ${\mathfrak{A}}^{\prime}$ over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-hard.
It is well-known that $K_{3}$ pp-constructs all finite relational structures (see, e.g, [3, Corollary 6.4.4]). Hence, by the transitivity of pp-constructability, every valued structure that pp-constructs $K_{3}$ pp-constructs all finite relational structures.
2.6. Fractional polymorphisms
@@
Let ${\mathfrak{A}}$ be a relational structure and $\ell\in\mathbb{N}$. An operation $f\in\operatorname{Pol}^{(\ell)}({\mathfrak{A}})$ is called a pseudo weak near unanimity (pwnu) polymorphism if there exist $e_{1},\dots,e_{\ell}\in\operatorname{End}({\mathfrak{A}})$ such that for every $x,y\in A$
$e_{1}f(y,x,\dots,x)=e_{2}f(x,y,x,\dots,x)=\dots=e_{\ell}f(x,\dots,x,y).$
We now introduce fractional polymorphisms of valued structures, which generalize polymorphisms of relational structures. Similarly as polymorphisms, fractional polymorphisms are an important tool for formulating tractability results and complexity classifications of VCSPs. For valued structures with a finite domain, our definition specialises to the established notion of a fractional polymorphism which has been used to study the complexity of VCSPs for valued structures over finite domains (see, e.g. [38]). Our definition is taken from [15] and allows arbitrary probability spaces in contrast to [39, 37, 40].
Definition 2.17 (fractional operation).
Let $\ell\in{\mathbb{N}}$. A fractional operation on $A$ of arity $\ell$ is a probability distribution
$\big{(}{\mathscr{O}}_{A}^{(\ell)},\mathcal{B}({\mathscr{O}}_{A}^{(\ell)}),\omega\colon\mathcal{B}({\mathscr{O}}_{A}^{(\ell)})\to[0,1]\big{)}.$
The set of all fractional operations on $A$ of arity $\ell$ is denoted by ${\mathscr{F}}^{(\ell)}_{A}$.
Definition 2.18.
@@
$E:=E_{\omega}[f\mapsto R(f(t^{1},\dots,t^{\ell}))]$
exists, and
$E\leq\frac{1}{\ell}\sum_{j=1}^{\ell}R(t^{j}).$ (2)
Note that (2) has the interpretation that the expected value of $R(f(t^{1},\dots,t^{\ell}))$ is at most the average of the values $R(t^{1}),\dots$, $R(t^{\ell})$.
Definition 2.19 (fractional polymorphism).
If a fractional operation $\omega$ improves every valued relation in a valued structure ${\mathfrak{A}}$, then $\omega$ is called a fractional polymorphism of ${\mathfrak{A}}$; the set of all fractional polymorphisms of ${\mathfrak{A}}$ is denoted by $\operatorname{fPol}({\mathfrak{A}})$.
Remark 2.20.
@@
Remark 2.22.
Let ${\mathfrak{A}}$ be a relational $\tau$-structure on the domain $A$. Then $\operatorname{Pol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}})$ (using the convention introduced above that an operation on $A$ can be viewed as a fractional operation). More concretely, for every $\ell\in\mathbb{N}$ we have that $\operatorname{fPol}^{(\ell)}({\mathfrak{A}})$ consists of precisely the fractional operations $\omega$ of arity $\ell$ such that $\omega(\operatorname{Pol}^{(\ell)}({\mathfrak{A}}))=1$. To see this, note that
$\mathscr{O}_{A}^{(\ell)}\setminus\operatorname{Pol}^{(\ell)}({\mathfrak{A}})=\bigcup_{t^{1},\dots,t^{\ell}\in R,s\in A^{k}\setminus R}\mathscr{S}_{(t^{1},\dots,t^{\ell}),s}$ (3)
and therefore a measurable set. Hence, $\operatorname{Pol}^{(\ell)}({\mathfrak{A}})$ is also measurable. Let $R\in\tau$ be of arity $k$ and let $t^{1},\dots,t^{\ell}\in A^{k}$. Note that by (1)
$E_{\omega}[f\mapsto R(f(t^{1},\dots,t^{\ell}))]=\sum_{s\in A^{k}}\omega(\mathscr{S}_{(t^{1},\dots,t^{\ell}),s})R(s).$
Therefore, if $\omega(\operatorname{Pol}^{(\ell)}({\mathfrak{A}}))=1$, either $R(t^{j})=\infty$ for some $j$ and (2) holds trivially, or $t^{1},\dots,t^{\ell}\in R$ and $\omega(\mathscr{S}_{(t^{1},\dots,t^{\ell}),s})=0$ whenever $s\notin R$. In this case (2) holds because both sides of the inequality are equal to 00. On the other hand, if $\omega(\mathscr{O}_{A}^{(\ell)}\setminus\operatorname{Pol}^{(\ell)}({\mathfrak{A}}))>0$, then by (3) there exist $t^{1},\dots,t^{\ell}\in R$ and $s\in A^{k}\setminus R$ such that $\omega(\mathscr{S}_{(t^{1},\dots,t^{\ell}),s})>0$. Then $E_{\omega}[f\mapsto R(f(t^{1},\dots,t^{\ell}))]=\infty$, which contradicts (2) since $R(t^{j})=0$ for all $j$.
Lemma 2.23 (Lemma 6.8 in [15]).
Let ${\mathfrak{A}}$ be a valued $\tau$-structure ${\mathfrak{A}}$ over a countable domain $A$. Then every valued relation $R\in\langle{\mathfrak{A}}\rangle$ is improved by all fractional polymorphisms of ${\mathfrak{A}}$.
Remark 2.24.
Recall that $\langle{\mathfrak{A}}\rangle=\langle\operatorname{Feas}({\mathfrak{A}})\rangle$ for essentially crisp valued structures ${\mathfrak{A}}$. Therefore, Lemma 2.23 implies that $\operatorname{fPol}({\mathfrak{A}})=\operatorname{fPol}(\operatorname{Feas}({\mathfrak{A}}))$.
2.7. Temporal valued structures
Let ${\mathfrak{A}}$ be a valued structure such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Then every $\tau$-expression $\phi$ (in particular, every valued relation of ${\mathfrak{A}}$) attains only finitely many values in ${\mathfrak{A}}$: if $\phi=\phi(x_{1},\dots,x_{k})$, we may have at most one value for every orbit of the action of $\operatorname{Aut}({\mathfrak{A}})$ on $\mathbb{Q}^{k}$. Since the group $\operatorname{Aut}(\mathbb{Q};<)$ has only finitely many orbits of $k$-tuples for every $k$, so does $\operatorname{Aut}({\mathfrak{A}})$. In particular, if $k=2$ and $a,b\in\mathbb{Q}$, $a<b$, then the values $\phi(a,a),\phi(a,b)$ and $\phi(b,a)$ do not depend on the choice of $a$ and $b$ and if $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};=)$, then $\phi(a,b)=\phi(b,a)$. As a consequence, for every $\tau$-expression $\phi(x_{1},\dots,x_{k},y_{1},\dots,y_{\ell})$ and $b\in A^{\ell}$ there exists $a^{*}\in A^{k}$ such that

### v2

@@ 2.5 Pp-constructions
Definition 7 (fractional map)
Let $A$ and $B$ be sets. A fractional map from $B$ to $A$ is a probability distribution $(A^{B},\mathcal{B}(A^{B}),\omega\colon\mathcal{B}(A^{B})\to[0,1]),$ that is, $\omega(A^{B})=1$ and $\omega$ is countably additive: if $S_{1},S_{2},\dots\in\mathcal{B}(A^{B})$ are disjoint, then
$\omega\left(\bigcup_{i\in{\mathbb{N}}}S_{i}\right)=\sum_{i\in{\mathbb{N}}}\omega(S_{i}).$
@@
The set $[0,1]$ carries the topology inherited from the standard topology on $\mathbb{R}$}. We also view $\mathbb{R}$$\cup\{\infty\}$ as a topological space with a basis of open sets given by all open intervals $(a,b)$ for $\in{\mathbb{R}}, a<ba<b$ and additionally all sets of the form $\{x\in{\mathbb{R}}\mid x>a\}\cup\{\infty\}$ (thus, $\mathbb{R}\cup\{\infty\}$ is equipped with its order topology when ordered in the natural way).
@@
Definition 8 (fractional homomorphism)
Let $\mathfrak{A}$} and $\mathfrak{B}$} be valued $\tau$-structures with domains $A$ and $B$, respectively. A fractional homomorphism from $\mathfrak{B}$} to $\mathfrak{A}$} is a  fractional map $\omega$ from $B$ to $A$ such that for every $R\in\tau$ of arity $k$ and every tuple $t\in B^{k}$ it holds for the random variable $X\colon A^{B}\rightarrow\mathbb{R}\cup\{\infty\}$ given by $f\mapsto R^{{\mathfrak{A}}}(f(t))$ that $E_{\omega}[X]$ exists and that $E_{\omega}[X]\leq R^{{\mathfrak{B}}}(t).$
We refer to [16] for a detailed introduction to fractional homomorphisms in full generality. We point out that fractional maps compose and the composition operation generalizes the standard composition of operations. Moreover, a composition of fractional homomorphisms is a fractional homomorphism [15, Proposition 5.9]. Observe that if $\mathfrak{A}$} is a countable valued structure and $R$ is a valued relation of $\mathfrak{A}$}, we have the following handy expression for $E_{\omega}[f\mapsto R(f(t))]$, which we sometimes use in the proofs (see [16] for a detailed derivation):
$E_{\omega}[f\mapsto R(f(t))]=\sum_{s\in A^{k}}R(s)\omega({\mathscr{S}}_{t,s}).$ (1)
All concrete fractional homomorphisms $\omega$ that appear in this paper are of a very special form, namely, there is a single $f\in A^{B}$ such that $\omega(\{f\})=1$. In this case, we also write $f$ instead of $\omega$. If $\omega$ is of this form, then for every $R\in\tau$ of arity $k$ and $t\in B^{k}$, the expected value in Definition 8 always exists and is equal to $R^{\mathfrak{A}}(f(t))$. If additionally $\mathfrak{A}$} and $\mathfrak{B}$} are crisp structures, then we call ffhomomorphism. It is easy to see that there are valued structures $\mathfrak{A}$} and $\mathfrak{B}$} with a fractional homomorphism from $\mathfrak{B}$} to $\mathfrak{A}$}, but no fractional homomorphism from $\mathfrak{B}$} to $\mathfrak{A}$} of the form $f\in A^{B}$ (see, e.g. [41, Example 2.6 and 3.15]).
Lemma 9
@@
Proof 1
Suppose that there exists a fractional homomorphism $\omega$ from $\mathfrak{B}$} to $\mathfrak{A}$}. Since $B$ is countable and $\operatorname{Aut}({\mathfrak{A}})\subseteq\operatorname{Aut}(\operatorname{Feas}({\mathfrak{A}}))$ is oligomorphic, it suffices to show that every finite substructure $\mathfrak{F}$} of $\operatorname{Feas}({\mathfrak{B}})$ has a homomorphism to $\operatorname{Feas}({\mathfrak{A}})$ (see, e.g., [17, Lemma 4.1.7]). Let $b_{1},\dots,b_{n}$ be the elements of $\mathfrak{F}$} and $b:=(b_{1},\dots,b_{n})$. By the (countable) additivity of probability distributions, there exists $a\in A^{n}$ such that $\omega(\mathscr{S}_{b,a})>0$. Let $f$ be the map that takes $b$ to $a$. Suppose that there exists $R\in\tau$ of arity $k$ and $t\in F^{k}$ such that $R^{{\mathfrak{A}}}(f(t))=\infty$. Since $\omega(\mathscr{S}_{b,a})>0$ and $\mathscr{S}_{b,a}\subseteq{\mathscr{S}}_{t,f(t)}$, we have $\omega({\mathscr{S}}_{t,f(t)})>0$, and thus $E_{\omega}[g\mapsto R^{{\mathfrak{A}}}(g(t))]=\infty$ by (1). Then $R^{{\mathfrak{B}}}(t)=\infty$, because $\omega$ is a fractional polymorphism. Hence, for every $R\in\tau$ of arity $k$ and $t\in F^{k}$ we have $R^{{\mathfrak{A}}}(f(t))<\infty$ whenever $R^{{\mathfrak{B}}}(t)<\infty$. Therefore, $f$ is a homomorphism from $\mathfrak{F}$} to $\operatorname{Feas}({\mathfrak{A}})$.
The final statement follows from the fact that every homomorphism is a fractional homomorphism and that ${\mathfrak{C}}=\operatorname{Feas}({\mathfrak{C}})$ for every crisp structure $\mathfrak{C}$}. $\qed$
We say that two valued $\tau$-structures $\mathfrak{A}$} and $\mathfrak{B}$} are fractionally homomorphically equivalent if there exists a fractional homomorphisms from $\mathfrak{A}$} to $\mathfrak{B}$} and from $\mathfrak{B}$} to $\mathfrak{A}$}. Since fractional homomorphisms compose [16], fractional homomorphic equivalence is indeed an equivalence relation on valued structures of the same signature.
Definition 10 (pp-construction)
@@
Remark 11
We remark that our definition of pp-constructability (Definition 10) applied to relational structures with oligomorphic automorphism groups coincides with the definition from [2]. Let $\mathfrak{A}$} and $\mathfrak{B}$} be relational structures, and suppose that $\mathfrak{A}$} $p$-constructs $\mathfrak{B}$}, that is, there is a pp-power $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} which is fractionally homomorphically equivalent to $\mathfrak{B}$}. Since $\mathfrak{A}$} is crisp, $\mathfrak{A}$$^{\prime}$ is essentially crisp and $\operatorname{Feas}({\mathfrak{A}}^{\prime})$ is a pp-power of $\mathfrak{A}$} with all relations primitively positively definable in $\mathfrak{A}$} when viewed over the domain $A$. By Lemma 9, $\operatorname{Feas}({\mathfrak{A}}^{\prime})$ is homomorphically equivalent to $\mathfrak{B}$}. This shows the claim.
The relation of pp-constructability is transitive: if $\mathfrak{A}$}, $\mathfrak{B}$}, and $\mathfrak{C}$} are valued structures such that $\mathfrak{A}$} $p$-constructs $\mathfrak{B}$} and $\mathfrak{B}$} $p$-constructs $\mathfrak{C}$}, then $\mathfrak{A}$} $p$-constructs $\mathfrak{C}$} [16, Lemma 5.14].
By $K_{3}$ we denote the complete graph on $3$ vertices. The following is a direct consequence of [16, Corollary 5.13] and [17, Corollary 6.7.13].
Lemma 12
Let $\mathfrak{A}$} be a valued structure such that $\operatorname{Aut}({\mathfrak{A}})$ is oligomorphic. If $\mathfrak{A}$} $p$-constructs $K_{3}$, then $\mathfrak{A}$} has a reduct $\mathfrak{A}$$^{\prime}$ over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-hard.
It is well-known that $K_{3}$ pp-constructs all finite relational structures (see, e.g, [17, Corollary 6.4.4]). Hence, by the transitivity of pp-constructability, every valued structure that pp-constructs $K_{3}$ pp-constructs all finite relational structures.
2.6 Fractional polymorphisms
@@
Let $\mathfrak{A}$} be a relational structure and $\ell\in\mathbb{N}$. An operation $f\in\operatorname{Pol}^{(\ell)}({\mathfrak{A}})$ is called a pseudo weak near unanimity (pwnu) polymorphism if there exist $e_{1},\dots,e_{\ell}\in\operatorname{End}({\mathfrak{A}})$ such that for every $x,y\in A$, $e_{1}f(y,x,\dots,x)=e_{2}f(x,y,x,\dots,x)=\dots=e_{\ell}f(x,\dots,x,y).$ We now introduce fractional polymorphisms of valued structures, which generalize polymorphisms of relational structures. Similarly to polymorphisms, fractional polymorphisms are an important tool for formulating tractability results and complexity classifications of VCSPs. For valued structures with a finite domain, our definition specialises to the established notion of a fractional polymorphism which has been used to study the complexity of VCSPs for valued structures over finite domains (see, e.g. [42]). Our definition is taken from [16] and allows arbitrary probability spaces in contrast to [44, 40, 43].
Definition 13 (fractional operation)
Let $\ell\in{\mathbb{N}}$. A fractional operation on $A$ of arity $\ell$ is a fractional map from $A^{\ell}$ to $A$. The set of all fractional operations on $A$ of arity $\ell$ is denoted by ${\mathscr{F}}^{(\ell)}_{A}$.
Definition 14
@@
$E:=E_{\omega}[f\mapsto R(f(t^{1},\dots,t^{\ell}))]\text{ exists, and }\quad E\leq\frac{1}{\ell}\sum_{j=1}^{\ell}R(t^{j}).$ (2)
Note that (2) has the interpretation that the expected value of $f\mapsto R(f(t^{1},\dots,t^{\ell}))$ is at most the average of the values $R(t^{1}),\dots$, $R(t^{\ell})$.
Definition 15 (fractional polymorphism)
A fractional operation which improves every valued relation in a valued structure $\mathfrak{A}$} is called a fractional polymorphism of $\mathfrak{A}$}; the set of all fractional polymorphisms of $\mathfrak{A}$} is denoted by $\operatorname{fPol}({\mathfrak{A}})$.
Remark 16
@@
Lemma 18 (Lemma 6.8 in [16])
Let $\mathfrak{A}$} be a valued $\tau$-structure $\mathfrak{A}$} over a countable domain $A$ and let $R$ be a valued relation over $A$. If $R\in\langle{\mathfrak{A}}\rangle$, then $R$ is improved by all fractional polymorphisms of $\mathfrak{A}$}.
Remark 19
If $\mathfrak{A}$} is a valued structure such that $\operatorname{Aut}({\mathfrak{A}})$ is oligomorphic, then the converse of Lemma 18 might be true as well; this is for instance the case if $\mathfrak{A}$} is a relational structure and $R$ a crisp relation [11].
Remark 20
Observe that $\operatorname{Pol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}})$ for every crisp structure $\mathfrak{A}$}. More concretely, for every $\ell\in\mathbb{N}$ we have that $\operatorname{fPol}^{(\ell)}({\mathfrak{A}})$ consists of precisely the fractional operations $\omega$ of arity $\ell$ such that $\omega(\operatorname{Pol}^{(\ell)}({\mathfrak{A}}))=1$. Also recall that, for essentially crisp valued structures $\mathfrak{A}$}, $\langle{\mathfrak{A}}\rangle=\langle\operatorname{Feas}({\mathfrak{A}})\rangle$. Therefore, Lemma 18 implies that $\operatorname{fPol}({\mathfrak{A}})=\operatorname{fPol}(\operatorname{Feas}({\mathfrak{A}}))$.
2.7 Temporal valued structures
Recall that a temporal valued structure is a a valued structure $\mathfrak{A}$} such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Every $\tau$-expression $\phi$ (in particular, every valued relation of $\mathfrak{A}$}) attains only finitely many values in $\mathfrak{A}$}: if $\phi=\phi(x_{1},\dots,x_{k})$, we may have at most one value for every orbit of the action of $\operatorname{Aut}({\mathfrak{A}})$ on $\mathbb{Q}^{k}$. Since the group $\operatorname{Aut}(\mathbb{Q};<)$ has only finitely many orbits of $k$-tuples for every $k$, so does $\operatorname{Aut}({\mathfrak{A}})$. In particular, if $k=2$ and $\in\mathbb{Q}, a<ba<b$, then the values $\phi(a,a)$, $\phi(a,b)$, and $\phi(b,a)$ do not depend on the choice of $a$ and $b$. If $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};=)$, then we additionally have $\phi(a,b)=\phi(b,a)$. As a consequence, for every $\tau$-expression $\phi(x_{1},\dots,x_{k},y_{1},\dots,y_{\ell})$ and $b\in A^{\ell}$ there exists $a^{*}\in A^{k}$ such that